40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 19 22:05:09 2016, Koji, Summary, IOO, PMC, IMC aligned. The ringdown PD/Lens removed. 
    Reply  Wed Dec 21 21:03:06 2016, Koji, Summary, IOO, IMC WFS tuning DSC_0142.JPG
       Reply  Thu Dec 22 18:39:09 2016, Koji, Summary, IOO, IMC WFS tuning WFS_error_noise.pdfWFS_error_noise_chans.pdf
          Reply  Fri Dec 23 20:53:44 2016, Koji, Summary, IOO, IMC WFS tuning DSC_0144.JPGDSC_0145.JPG
             Reply  Fri Dec 23 21:05:56 2016, Koji, Summary, IOO, IMC WFS tuning IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfIMC_WFS_161221_table1.pdfIMC_WFS_161221_table2.pdfIMC_WFS_161221.xlsx.zip
                Reply  Sun Dec 25 14:39:59 2016, Koji, Summary, IOO, IMC WFS tuning 
                   Reply  Mon Dec 26 12:45:31 2016, Koji, Summary, IOO, IMC WFS tuning IMC_REFL_TRANS_26hrs.pngIMC_TRANS_P_Y_26hrs.png
          Reply  Thu Dec 29 13:22:21 2016, rana, Summary, IOO, IMC WFS tuning 
             Reply  Thu Dec 29 16:52:51 2016, Koji, Summary, IOO, IMC WFS tuning 
                Reply  Thu Dec 29 21:35:30 2016, rana, Summary, IOO, IMC WFS tuning Screen_Shot_2017-01-03_at_5.55.13_PM.png
                   Reply  Tue Jan 24 01:04:16 2017, gautam, Summary, IOO, IMC WFS RF power levels MC1_misalignment.png
                      Reply  Fri Jan 27 00:14:02 2017, gautam, Summary, IOO, IMC WFS RF power levels 
Message ID: 12680     Entry time: Wed Dec 21 21:03:06 2016     In reply to: 12679     Reply to this: 12682
Author: Koji 
Type: Summary 
Category: IOO 
Subject: IMC WFS tuning 

- Updated the circuit diagrams:

IMC WFS Demodulator Board, Rev. 40m https://dcc.ligo.org/LIGO-D1600503

IMC WFS Whitening Board, Rev. 40m https://dcc.ligo.org/LIGO-D1600504

- Measured the noise levels of the whitening board, demodboard, and nominal free running WFS signals.

- IMC WFS demod phases for 8ch adjusted

Injected an IMC PDH error point offset (@1kHz, 10mV, 10dB gain) and adjusted the phase to have no signal in the Q phase signals.

- The WFS2 PITCH/YAW matrix was fixed

It was found that the WFS heads were rotated by 45 deg (->OK) in CW and CCW for WFS1 and 2, respectively (oh!), while the input matrices were identical! This made the pitch and yaw swapped for WFS2. (See attachment)

- Measured the TFs MC1/2/3 P/Y actuation to the error signals

Attachment 1: DSC_0142.JPG  204 kB  | Hide | Hide all
DSC_0142.JPG
ELOG V3.1.3-